A ball is dropped from a height of 20 feet and bounces back to 60% of the previous height. What is the total distance traveled?
Answer by ikleyn(52787)

.
A ball is dropped from a height of 20 feet and bounces back to 60% of the previous height. What is the total distance traveled?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

20 + 0.6*20 = 20 + 12 = 32 feet.